OpenStudy (cwtan):

Given A,B,C are the angles of an triangle. Shows that sin (A)+sin(B)+sin(C)=4 cos (A/2) cos (B/2) cos (C/2)

OpenStudy (anonymous):

A+B+C= Pi then taking B+C=D so sinA=Π−D=cosD=cos(B+C) .... cos(B+C) + sinB+ sinC =cosBcosC-sinBsinC + sinB + sinC

OpenStudy (cwtan):

Erm if it includes angle Please show it in degree... Radian a bit confusing me... Tq

OpenStudy (anonymous):

Let \[ x + y + z =\frac \pi 2\\ \sin (2 x)+\sin (2 y)+\sin (2 z)=2 \sin (x+y) \cos (x-y)+2 \sin (z) \cos (z)=\\2 \sin (x+y) \cos (x-y)+2 \sin \left(\frac{\pi }{2}-(x+y)\right) \cos \left(\frac{\pi }{2}-(x+y)\right)\\=2 \sin (x+y) \cos (x-y)+2 \sin (x+y) \cos (x+y)\\=2 \sin (x+y) (\cos (x-y)+\cos (x+y))=\\2 \sin (x+y) (2 \cos (x) \cos (y))=\\4 \cos (x) \cos (y) \sin \left(\frac{\pi }{2}-z\right)=\\4 \cos (x) \cos (y) \cos (z) \] This is the solution of your problem.

OpenStudy (anonymous):

@eliassaab but u have to prove as 4 cos (A/2) cos (B/2) cos (C/2)

OpenStudy (anonymous):

A=2x, x =A/2 B= 2y, y=B/2 C= 2y, y=C/2 Got it.